OpenStudy (dtan5457):

A small-appliance manufacturer finds that the profit p (in dollars) generated by producing x microwave ovens per week is given by the formula P=1/10x(300-x) produced that 0<=x<=200. How many ovens must be manufactured in a given week to generate a profit of $1250?

OpenStudy (dtan5457):

Anyone have a clue?

OpenStudy (anonymous):

2105.60 = x/10 * (300 - x) 300x - x² = 21056 x² - 300x + 21056 = 0 x = (300 +/- √(90,000 - 4(21056))/2 = (300 +/- 76)/2 x = 112 or x = 188

OpenStudy (dtan5457):

where'd you get the 2105.60?

OpenStudy (dtan5457):

Anyone with a more worded answer so I can understand better?